OpenStudy (anonymous):

Find the intercepts for y = x − 4

OpenStudy (anonymous):

There are two of them. For the y-intercept, let x = 0 then solve as shown below y = x - 4 y = 0 - 4 y intercept = -4 Can you do the same for the x-intercept?

OpenStudy (anonymous):

is it 1 ?

OpenStudy (anonymous):

lets see For the x-intercept, let y = 0 then solve as shown below 0 = x - 4 x = 4 The x-intercept is 4 and the y-intercept is -4.